What volume would be occupied by 3.5 mol of oxygen gas at a pressure of 768 mmHg and a temperature of 25 degrees Celsius?
Answer #1

Given:

P = 768.0 mm Hg

= (768.0/760) atm

= 1.0105 atm

n = 3.5 mol

T = 25.0 oC

= (25.0+273) K

= 298 K

use:

P * V = n*R*T

1.0105 atm * V = 3.5 mol* 0.08206 atm.L/mol.K * 298 K

V = 84.697 L

Answer: 84.7 L